OpenStudy (anonymous):

list five elements of V, where V is the set of all real real valued differentiable functions whose domain is the set of all real numbers and having the property that f'(x)=0

OpenStudy (anonymous):

y=1, y=2, y=3...?

OpenStudy (anonymous):

sorry, i have to fix.. f'(1)=0

OpenStudy (anonymous):

so just restricting ourselves to f(x) = ax^2 + bx + c = 0, find solutions to 2a + b = 0

OpenStudy (jamesj):

f(x) = (x-1)^2 is one f(x) = (x-1)^3 is another f(x) = (x-1)^4 another ...

OpenStudy (anonymous):

yeaa..